OpenStudy (anonymous):

4. Evaluate the expression for a = –1 and b = 5. (ab)2

OpenStudy (anonymous):

If a = -1 b = 5 Then (ab)2 =(-1 x 5) =-5

OpenStudy (anonymous):

Sorry =(-5)2 =-25

OpenStudy (anonymous):

this is what i did ........... (ab)2=25

OpenStudy (anonymous):

But a is a -1

OpenStudy (anonymous):

yup (ab)2 = to -25 thank u though !

OpenStudy (anonymous):

ya welcome

OpenStudy (anonymous):

wanna help me with one more?

OpenStudy (anonymous):

Simplify the expression. 9xy2 – 11xy2=

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):

k ive got plenty lol lets try and do them together if u dont mind ?

OpenStudy (anonymous):

9 xy2-11 xy2= =9(xy)2- 11(xy)2 -xy2=9-11 xy2=-9+11 xy2=2 xy=2/2 xy=1?

OpenStudy (anonymous):

dont ask me lol, is that right ?

OpenStudy (anonymous):

-2xy2

OpenStudy (anonymous):

dont solve it lol just simplify it

OpenStudy (anonymous):

ok